Ingrid invested $24,000 in two separate accounts, a mutual fund and a CD that pay 8% and 9% annual interest, respectively. If her annual interest was $2,020 how much did she invest in each account?
Answer by stanbon(75887)

Ingrid invested $24,000 in two separate accounts, a mutual fund and a CD that pay 8% and 9% annual interest, respectively. If her annual interest was $2,020 how much did she invest in each account?
-----
Quantity Eq:: e + n = 24000
Interest Eq: 0.08e + 0.09n = 2020
--------------------
Modify for elimination:
8e + 8n = 8*24000
8e + 9n = 202000
---------------------------------
Subtract and solve for "n"::
n = $10,000 (amt invested at 9%)
---
e = 24000-n = $14000 (amt invested at 8%)
----------
Cheers,
Stan H.
-----------